OpenStudy (anonymous):

find the equation with the given slope and containing the given point. Slope 3: through (1,2)

OpenStudy (anonymous):

you wanna use this: y-y1=m(x-x1) y-2=3(x-1) y-2=3x-3 y=3x-1

OpenStudy (anonymous):

y-2=3(x-1)*

OpenStudy (katrinakaif):

Plug it into this equation: (y-y1) = m(x-x1) As (x1,y1) = (1,2) and m = 3 as its slope.. =)

OpenStudy (anonymous):

the answer says f(x) = 3x-1

OpenStudy (anonymous):

but it only has 2 points not 4

OpenStudy (anonymous):

the y and the x are ur variables the y1 and x1 are the points they gave u

OpenStudy (anonymous):

so y2 and x2 would be just zero?

OpenStudy (anonymous):

there is no y2 or x2

OpenStudy (anonymous):

so how did they get the answer

OpenStudy (anonymous):

y-y1=m(x-x1) m is the slope (y1,x1) is the point so m=3 y1=2 x1=1 so y-2=3(x-1) y-2=3x-3 y-2+2=3x-3+2 y=3x-1

OpenStudy (anonymous):

sorry its (x1,y1) not (y1,x1)

OpenStudy (anonymous):

oh ok thank u